40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3991     Entry time: Mon Nov 29 22:50:07 2010
Author: Suresh 
Type: Update 
Category: SUS 
Subject: ETMU05 Side Magnets glued back  

[Suresh, Jenne]

ETMU05 : Gluing Side magnets back on to the optic.

The following steps taken in this process:

1) The two magnet+dumbell units which had come loose from the optic needed to be cleaned.  A lint free wipe was placed on the table top and a few cc of acetone was poured on to it.  The free end of the dumbbell was then scrubbed on this wipe till the surface regained its shine.  The dumbell was held at its narrow part with a forceps to avoid any strain on the magnet-dumbbell joint.

2) The optic was then removed from its gluing fixture (by loosening only one of the three retaining screws) and placed in an Al ring. The glue left behind by the side magnets was scrubbed off with a optical tissue wetted with Acetone. 

3) The optic was returned to the gluing fixture.  The position of the optic was checked by inserting the brass portion of the gripper and making sure that the face magnets are centered in it [Jenne doubled checked to be sure we got everything right].

4) The side magnets were glued on and the optic in the fixture has been placed in the foil-house.

If all goes well we will be able to balance the ETMU05 and give it to Bob for baking.

 

ETMU07 : It is still in the oven and we need to ask Bob to take out. It will be available for installation in the 40m tomorrow.  

 

ELOG V3.1.3-